OpenStudy (anonymous):

Predict the y-value if x = -3. x: -1, 0, 1, 2. y: 5, 3, 1, -1. I thought was the x value was: -1, 0, 1, 2? Not x=-3.... how can this be? ?????? Please help :(

OpenStudy (amistre64):

you have a constant difference between y values, can you spot that difference?

OpenStudy (anonymous):

the difference in x's pattern is adding 1 from -1. in y's pattern it's by adding -2 from 5. is that what you were asking?

OpenStudy (amistre64):

yes, thats what I was asking; since y changes by -2 when x changes by +1 we can create a slope by stacking y/x slope = -2/1, or simply -2 notice that when x=0 we have a y value, known as the y intercept value; do you see it?

OpenStudy (amistre64):

this method is a bit overkill for this problem, but its a good idea to have

OpenStudy (anonymous):

I think I understand it better :)

OpenStudy (amistre64):

using a slope intercept from of a line: y = mx+b we can construct a model to predict the unknown value y = -2x +3

OpenStudy (amistre64):

otherwise, we could have simply went backwards from 5 to get 2 extra spots -3 -2 -1 9 7 5 +2 +2

OpenStudy (anonymous):

Okay thank you so much :) !!